1. A drag racer constantly accelerates at a rate of 39 m/s2

to increase its speed in a short amount of time. If the drag racer started at rest, what distance did it have to travel to reach a velocity of 33 m/s, E?




1
Expert's answer
2021-10-13T15:41:01-0400
332=2(39)ss=14 m33^2=2(39)s\\s=14\ m